Solve 3^2x-1 = 5^x for x
Answer by stanbon(75887)

Solve 3^2x-1 = 5^x for x
-----------------
(2x-1)*log(3) = x*log(5)
------
(2x*log(3)-log(3) - [x*log(5)] = 0
-----
x[2log(3) - (log(5)] = log(3)
-------
x[log(9)-log(5)] = log(3)
-------
x = 0.4771/0.2553
x = 1.8690
---
Cheers,
Stan H.
----------